# A Principle of Targeted Intervention for Multi-Agent Reinforcement Learning

### v3

@@ H.2 Implementation of Baselines
Base MARL. This refers to the standard, unmodified (“vanilla”) version of the underlying MARL algorithm (e.g., IQL, VDN, IPPO). It uses the same network architecture and hyperparameters as our method but excludes both the pre-policy module and the additional intrinsic reward to maximize.
@@
Global Intervention. This approach represents a category of algorithms implementing the global intervention paradigm. In our experiments, we follow the specific implementation of two representative methods for this approach: LIIR [8] for policy-based algorithms, and LAIES [29] for value-based algorithms. Both methods in principle allocate an additional intrinsic reward to all agents to maximize. This contrasts the Intrinsic Reward approach above, where only one targeted agent is provided with an intrinsic reward to maximize.
@@ H.3 GNN Implementation
The foundation of our targeted intervention method (Pre-Strategy Intervention) is formed by the s-reachability graph criterion, introduced to identify the relevant policies of other agents for decision-making [17]. However, considering only policies is insufficient, as an agent’s policy may depend on other elements within the game [80]. Algorithm 1 provides an implementation how we can build connection with a causal graph structure for pre-policy learning. By leveraging this graph structure, we incorporate prior knowledge about the game to help guide agents’ policy-making in practical. The feasibility of learning the causal graph during training has been demonstrated by Richens and Everitt [89], where agents can learn the causal model implicitly during interaction with the environment.
Architecture. If the adjacency matrix is not predefined, the GNN processes the observation vectors by first encoding them into logits, which are used to generate a soft adjacency matrix via the gumbel-softmax technique [90]. This matrix defines the relationship between features in the observations. Once the adjacency matrix is formed, a graph convolutional layer applies message passing to update the features of each node based on its neighbors [84]. The output node features are then aggregated using a mean-pooling operation to produce a graph embedding. This embedding is used for further processing or decision-making.
@@ H.4 MPE
Multi-Agent Particle Environment (MPE). In the MPE Simple Spread task [19], our setup features 3 agents in a 2D continuous space cooperatively navigating to 3 landmarks while avoiding collisions. We apply our targeted intervention to one agent, using an intrinsic reward defined by the negative distance to its selected landmark. This is a simultaneous-move environment.
@@ H.4.1 Intrinsic reward
Scenario 2 (Dynamic Target Landmark - Farthest from Teammates). The targeted agent receives an intrinsic reward for approaching the landmark currently farthest from its other two teammates. This encourages strategic positioning relative to teammates. The farthest landmark $\mathbf{l}^{*}$ is determined dynamically at each step as follows:
@@ H.5 Hanabi
Hanabi. Team success in the Hanabi card game heavily relies on shared conventions—implicit rules that enhance predictability and coordination [20]. Our experiments study the 2-agent version of the game. We investigate if our targeted intervention method (Pre-Strategy Intervention), applied to a single agent, can effectively promote adherence to such conventions and thereby improve overall team coordination. Specifically, we guide the targeted agent towards one of two common Hanabi conventions as a additional desired outcome: “5 Save” (prioritizing hints for unique, high-value cards to prevent accidental discards) or “The Chop” (making discards predictable by targeting the newest, unhinted card).
@@ H.5.2 Intrinsic Reward
Convention 1 (5 Save): This convention is crucial for maximizing potential scores, as rank 5 cards are unique for each suit and essential for completing a firework. It aims to enhance team coordination by guiding agents to provide timely and clear hints about rank 5 cards held by teammates. This is especially important if those cards are currently unhinted and therefore at high risk of being accidentally discarded by a teammate unaware of their value. The objective is to ensure these vital cards are identified, preserved, and eventually played successfully. The specific logic for calculating the intrinsic reward, which encourages the targeted agent’s convention-adherent hinting actions (e.g., when a teammate holds an unhinted rank 5 card), is detailed in Algorithm 2.
Convention 2 (The Chop:) This convention promotes predictability in discarding actions, which in turn helps improve team efficiency and reduce miscoordination. It establishes a clear default discard target, often targeted as the newest card in a player’s hand that has not yet received any hint (the “chop” card). By adhering to this rule, especially when no other plays or hints are clearly beneficial, an agent provides a predictable pattern for discards. This allows teammates to make more informed inferences about the game state and their own hands. The intrinsic reward encourages the targeted agent to follow this specific discard protocol. The detailed logic for identifying the “chop” card and calculating the reward based on the agent’s discard action is provided in Algorithm 3.
@@ I.1.1 Detailed Results of the Scenario shown in the Main Paper
Figure 9: Detailed results of the scenario of MPE shown in the main body of paper. For each base MARL algorithm (IQL, VDN, and QMIX), extrinsic and intrinsic returns are displayed side by side. The algorithm implemented in parameter sharing is denoted as PS, while that implemented in non-parameter sharing is denoted as NP.
@@ I.2.1 Detailed Results of the Scenario shown in the Main Paper
We now present detailed results for the Hanabi scenario discussed in the main paper (using the “5 Save” convention as the additional desired outcome), comparing our targeted intervention approach (Pre-Strategy Intervention) against base MARL methods. These comparisons are conducted across various underlying MARL algorithms, with results shown in Figure 11. The baseline methods are implemented in two common settings: parameter sharing (PS) and non-parameter sharing (NP). Our Pre-Strategy Intervention is implemented using a non-parameter sharing architecture, reflecting more realistic scenarios where agents may operate with decentralized learning and not directly exchange parameters during training.
@@
Figure 11: Detailed results of the scenario of Hanabi shown in the main body of paper (with the convention as “5 Save”). For each base MARL algorithm (IPPO, MAPPO, PQN-IQL and PQN-VDN), extrinsic and intrinsic returns are displayed side by side. The algorithm implemented in parameter sharing is denoted as PS, while that implemented in non-parameter sharing is denoted as NP. Our method is implemented in non-parameter sharing.
@@ I.2.2 Detailed Results of an Additional Scenario
Figure 12: Detailed results of the scenario of Hanabi shown in the main body of paper (with the convention as “The Chop”). For each base MARL algorithm (IPPO, MAPPO, PQN-IQL and PQN-VDN), extrinsic and intrinsic returns are displayed side by side. The algorithm implemented in parameter sharing is denoted as PS, while that implemented in non-parameter sharing is denoted as NP. Our method is implemented in non-parameter sharing.
@@ I.3 Additional Results Analysis
Generalizability Across Different Additional Desired Outcomes and Base MARL Algorithms. As shown in Figures 9, 10, 11, and 12, our targeted intervention approach (Pre-Strategy Intervention) consistently outperforms baseline approaches across various base MARL algorithms. Additionally, the results across different experimental scenarios (each with distinct additional desired outcomes, such as specific landmark targeting in MPE or convention adherence in Hanabi) suggest that incorporating well-defined additional desired outcomes via the Pre-Strategy Intervention can effectively promote the agent learning process towards improved overall task performance.
@@
I.4 Additional Comparison with Global Intervention Paradigm in Hanabi
Figure 13: Comparison of extrinsic returns in Hanabi, evaluating our Pre-Strategy Intervention (PSI) implementing the targeted intervention paradigm against an approach implementing the global intervention paradigm (i.e. applying intrinsic rewards to all agents [29]) across various MARL algorithm backbones.
Figure 13 provides further evidence in the Hanabi environment, specifically utilizing value-based MARL algorithm backbones (PQN-VDN and PQN-IQL), for the comparison between our Pre-Strategy Intervention implementing the targeted intervention paradigm and an approach implementing the global intervention paradigm. The results for extrinsic returns (task completion) are consistent with the findings presented in the main paper (e.g., Figure 5(d), which used different backbones). Our targeted intervention approach, when applied with either PQN-VDN or PQN-IQL, generally leads to higher final performance compared to the chosen global intervention approach (LAIES [29]) where all agents are simultaneously guided.
@@ I.5 Experiment with Noisy Observation
I.5.1 Both training and testing contain noise
In this scenario, the uniform noise (scaled by 0.2) perturbs the card belief distributions during both training and testing, simulating persistent observation uncertainty. As shown in Figures 14(a) and 14(b), our method still generally outperforms the base MARL algorithms in primary task completion (extrinsic return). Furthermore, the adherence to the additional desired outcome (intrinsic return) remains effective, indicating robustness when agents can adapt to consistent noise during training.
I.5.2 Only training environments contain noise
In this scenario, the noise (scaled by 0.05) is present only during training, simulating a scenario where the deployment environment has less noise than the training conditions. Figures 14(c) and 14(d) show that our method maintains an advantage in task completion over the baselines. However, the performance regarding the additional desired outcome (intrinsic return) is notably impaired compared to the noise-free or consistently noisy cases, performing similarly to the non-parameter sharing baseline.
I.5.3 Only testing environments contain noise
This scenario introduces the noise (scaled by 0.02) only during testing, simulating deployment in an environment with unexpected observation noise. As seen in Figures 14(e) and 14(f), the performance of our method aligns closely with the parameter-sharing baseline in terms of extrinsic return (with overlapping confidence intervals). Similarly, there is no significant advantage in achieving the additional desired outcome (intrinsic return) compared to the non-parameter sharing baseline.
@@ I.6 Hanabi 4 Players
Figure 15: Performance comparison in 4-player Hanabi, with PQN-VDN as the base algorithm for all methods. Our Pre-Strategy Intervention (PSI) is compared against the Base MARL, the Global Intervention and the Intrinsic Reward baselines.
To further validate the scalability of our Pre-Strategy Intervention, we conducted new experiments on the 4-player version of the Hanabi challenge. This setting serves as a more demanding benchmark for coordination, as the strategic complexity and communication burden are known to increase dramatically with the number of players [91]. The results presented in Figure 15 demonstrate that our method achieves significantly higher scores than both the base MARL algorithms and the approach of the global intervention paradigm. This strong performance in a larger-scale setting provides robust evidence of our method’s ability to effectively handle more complex coordination challenges.
@@ I.7 Heterogeneous MPE
To evaluate the performance of our targeted intervention approach (Pre-Strategy Intervention) in settings with diverse agent capabilities, we designed a heterogeneous MPE scenario. This environment features three agents with asymmetric speeds: one “sprinter” agent with five times the normal speed and two normal-speed agents, with all methods built upon the IQL algorithm. The results, shown in Figure 16, reveal a consistent trend: intervening on a normal-speed agent leads to a noticeable improvement in final team performance (extrinsic return) compared to intervening on the uniquely capable sprinter, although the magnitude of this difference is modest in this specific environment. Our analysis suggests this is because the system’s primary limitation: the coordination bottleneck between the two normal-speed agents, when intervening on the sprinter agent. Even a subtle increase in the predictability of one normal agent’s behavior through our intervention can alleviate this core challenge, leading to a more effective overall team strategy and demonstrating our approach’s potential in heterogeneous teams where targeted coordination adjustments can yield benefits.
The above results and analysis suggest that further investigation is needed to determine which agent to intervene on in future work.
@@
Our work on Pre-Strategy Intervention (implementing the targeted intervention paradigm) in multi-agent reinforcement learning (MARL) introduces a theoretical framework for agent coordination. As foundational research conducted entirely in simulated environments, the immediate societal impact is limited. However, we acknowledge that theoretical advances often lay groundwork for future applications. If extended to practical domains, this research could potentially enable more efficient coordination in multi-agent systems by focusing guidance on specific agents rather than requiring direct control of an entire system. Our targeted intervention paradigm could reduce computational and communication overhead in domains like robotic teams, traffic management, or resource allocation. At the same time, we recognize that techniques for influencing multi-agent systems through targeted intervention could potentially be misapplied if deployed in socially sensitive contexts without appropriate oversight. Future work that moves this theoretical foundation toward real-world applications should carefully consider application-specific ethical implications and incorporate appropriate safeguards against potential misuse. As our current work remains at the algorithmic level with experiments limited to standard MARL benchmarks (MPE and Hanabi), we do not anticipate immediate societal consequences from this research contribution.

### v4

@@ H.2 Implementation of Baselines
Base MARL. This refers to the standard, unmodified (“vanilla”) version of the underlying MARL algorithm. For MPE, these include IQL [tan1993multi], VDN [sunehag2017value] and QMIX [rashid2020monotonic]. For Hanabi, we use IPPO [schulman2017proximal], MAPPO [yu2022surprising], and variants of PQN [gallici2024simplifying], specifically PQN-IQL and PQN-VDN. All Base MARL variants use the same network architecture and hyperparameters as our method but exclude both the pre-policy module and the additional intrinsic reward to maximize.
@@
Global Intervention. This approach represents a category of algorithms implementing the global intervention paradigm. In our experiments, we follow the specific implementation of two representative methods for this approach: LIIR [du2019liir] for policy-based algorithms, and LAIES [liu2023lazy] for value-based algorithms. Both methods in principle allocate an additional intrinsic reward to each agent. This contrasts the Intrinsic Reward approach above, where only one targeted agent is provided with an intrinsic reward to maximize. For LAIES, the intrinsic reward is implemented by a function calculating the causal effect of their actions on external states. The LAIES framework proposes two types of intrinsic motivation: Individual Diligence (IDI) and Collaborative Diligence (CDI). Notably, the training process for the CDI component relies on access to the joint actions of all agents. As our proposed method and the other baselines in our experiments operate without leveraging joint action information during training, we implemented the LAIES baseline using solely its IDI component to maintain consistency and ensure a fair comparison.
@@ H.3 GNN Implementation
The foundation of our targeted intervention method (Pre-Strategy Intervention) is formed by the s-reachability graph criterion, introduced to identify the relevant policies of other agents for decision-making [koller2003multi]. However, considering only policies is insufficient, as an agent’s policy may depend on other elements within the game [hammond2023reasoning]. Algorithm 1 provides an implementation how we can build connection with a causal graph structure for pre-policy learning. By leveraging this graph structure, we incorporate prior knowledge about the game to help guide agents’ policy-making in practical. The feasibility of learning the causal graph during training has been demonstrated by Richens and Everitt [richens2024robust], where agents can learn the causal model implicitly during interaction with the environment.
Architecture. If the adjacency matrix is not predefined, the GNN processes the observation vectors by first encoding them into logits, which are used to generate a soft adjacency matrix via the gumbel-softmax technique [jang2016categorical]. This matrix defines the relationship between features in the observations. Once the adjacency matrix is formed, a graph convolutional layer applies message passing to update the features of each node based on its neighbors [pearl2014probabilistic]. The output node features are then aggregated using a mean-pooling operation to produce a graph embedding. This embedding is used for further processing or decision-making.
@@ H.4 MPE
In the MPE Simple Spread task [lowe2017multi], our setup features 3 agents in a 2D continuous space cooperatively navigating to 3 landmarks while avoiding collisions. We apply our targeted intervention to one agent, using an intrinsic reward defined by the negative distance to its selected landmark. This is a simultaneous-move environment.
@@ H.4.1 Intrinsic reward
Scenario 2 (Dynamic Target Landmark — Farthest from Teammates). The targeted agent receives an intrinsic reward for approaching the landmark currently farthest from its other two teammates. This encourages strategic positioning relative to teammates. The farthest landmark $\mathbf{l}^{*}$ is determined dynamically at each step as follows:
@@ H.5 Hanabi
Team success in the Hanabi card game heavily relies on shared conventions—implicit rules that enhance predictability and coordination [Bard_2020]. Our experiments study the 2-agent version of the game. We investigate if our targeted intervention method (Pre-Strategy Intervention), applied to a single agent, can effectively promote adherence to such conventions and thereby improve overall team coordination. Specifically, we guide the targeted agent towards one of two common Hanabi conventions as a additional desired outcome: “5 Save” (prioritizing hints for unique, high-value cards to prevent accidental discards) or “The Chop” (making discards predictable by targeting the newest, unhinted card).
@@ H.5.2 Intrinsic Reward
Convention 1 (5 Save). This convention is crucial for maximizing potential scores, as rank 5 cards are unique for each suit and essential for completing a firework. It aims to enhance team coordination by guiding agents to provide timely and clear hints about rank 5 cards held by teammates. This is especially important if those cards are currently unhinted and therefore at high risk of being accidentally discarded by a teammate unaware of their value. The objective is to ensure these vital cards are identified, preserved, and eventually played successfully. The specific logic for calculating the intrinsic reward, which encourages the targeted agent’s convention-adherent hinting actions (e.g., when a teammate holds an unhinted rank 5 card), is detailed in Algorithm 2.
Convention 2 (The Chop). This convention promotes predictability in discarding actions, which in turn helps improve team efficiency and reduce miscoordination. It establishes a clear default discard target, often targeted as the newest card in a player’s hand that has not yet received any hint (the “chop” card). By adhering to this rule, especially when no other plays or hints are clearly beneficial, an agent provides a predictable pattern for discards. This allows teammates to make more informed inferences about the game state and their own hands. The intrinsic reward encourages the targeted agent to follow this specific discard protocol. The detailed logic for identifying the “chop” card and calculating the reward based on the agent’s discard action is provided in Algorithm 3.
@@ I.1.1 Detailed Results of the Scenario shown in the Main Paper
Figure 8: Detailed results of the scenario of MPE shown in the main paper. For each base MARL algorithm (IQL, VDN, and QMIX), extrinsic and intrinsic returns are displayed side by side. The algorithm implemented in parameter sharing is denoted as PS, while that implemented in non-parameter sharing is denoted as NP.
@@ I.2.1 Detailed Results of the Scenario shown in the Main Paper
We now present detailed results for the Hanabi scenario discussed in the main paper (using the “5 Save” convention as the additional desired outcome), comparing our targeted intervention approach (Pre-Strategy Intervention) against base MARL methods. These comparisons are conducted across various underlying MARL algorithms, with results shown in Figure 10. The baseline methods are implemented in two common settings: parameter sharing (PS) and non-parameter sharing (NP). Our Pre-Strategy Intervention is implemented using a non-parameter sharing architecture, reflecting more realistic scenarios where agents may operate with decentralized training and not directly exchange parameters during training.
@@
Figure 10: Detailed results of the scenario of Hanabi shown in the main paper (with the convention as “5 Save”). For each base MARL algorithm (IPPO, MAPPO, PQN-IQL and PQN-VDN), extrinsic and intrinsic returns are displayed side by side. The algorithm implemented in parameter sharing is denoted as PS, while that implemented in non-parameter sharing is denoted as NP. Our method is implemented in non-parameter sharing.
@@ I.2.2 Detailed Results of an Additional Scenario
Figure 11: Detailed results of the scenario of Hanabi shown in the main paper (with the convention as “The Chop”). For each base MARL algorithm (IPPO, MAPPO, PQN-IQL and PQN-VDN), extrinsic and intrinsic returns are displayed side by side. The algorithm implemented in parameter sharing is denoted as PS, while that implemented in non-parameter sharing is denoted as NP. Our method is implemented in non-parameter sharing.
@@ I.3 Additional Results Analysis
Generalizability Across Different Additional Desired Outcomes and Base MARL Algorithms. As shown in Figures 8, 9, 10, and 11, our targeted intervention approach (Pre-Strategy Intervention) consistently outperforms baseline approaches across various base MARL algorithms. Additionally, the results across different experimental scenarios (each with distinct additional desired outcomes, such as specific landmark targeting in MPE or convention adherence in Hanabi) suggest that incorporating well-defined additional desired outcomes via the Pre-Strategy Intervention can effectively promote the agent learning process towards improved task completion performance.
@@
I.4 Additional Comparison with Additional-Outcome-Free Approaches in Hanabi
Figure 12: Comparison of extrinsic returns in Hanabi, evaluating our Pre-Strategy Intervention (PSI) implementing the targeted intervention paradigm against LAIES [liu2023lazy]) across various MARL algorithm backbones.
Figure 12 provides further evidence in the Hanabi environment, specifically utilizing value-based MARL algorithm backbones (PQN-VDN and PQN-IQL), for the comparison between our Pre-Strategy Intervention (PSI) implementing the targeted intervention paradigm and the LAIES [liu2023lazy] implementing the global intervention paradigm that only considers task completion. The results for extrinsic returns (task completion) are consistent with the findings presented in the main paper (e.g., Figure 5(d), which used different backbones). When applied with either PQN-VDN or PQN-IQL, PSI generally leads to higher performance than LAIES.
@@ I.5 Experiment with Noisy Observation
I.5.1 Both Training and Testing Containing Noise
In this scenario, the uniform noise (scaled by 0.2) perturbs the card belief distributions during both training and testing, simulating persistent observation uncertainty. As shown in Figures 13(a) and 13(b), our method still generally outperforms the base MARL algorithms in primary task completion (extrinsic return). Furthermore, the adherence to the additional desired outcome (intrinsic return) remains effective, indicating robustness when agents can adapt to consistent noise during training.
I.5.2 Only Training Environments Containing Noise
In this scenario, the noise (scaled by 0.05) is present only during training, simulating a scenario where the deployment environment has less noise than the training conditions. Figures 13(c) and 13(d) show that our method maintains an advantage in task completion over the baselines. However, the performance regarding the additional desired outcome (intrinsic return) is notably impaired compared to the noise-free or consistently noisy cases, performing similarly to the non-parameter sharing baseline.
I.5.3 Only Testing Environments Containing Noise
This scenario introduces the noise (scaled by 0.02) only during testing, simulating deployment in an environment with unexpected observation noise. As seen in Figures 13(e) and 13(f), the performance of our method aligns closely with the parameter-sharing baseline in terms of extrinsic return (with overlapping confidence intervals). Similarly, there is no significant advantage in achieving the additional desired outcome (intrinsic return) compared to the non-parameter sharing baseline.
@@ I.6 Hanabi 4 Players
Figure 14: Performance comparison in 4-player Hanabi, with PQN-VDN as the base algorithm for all methods. Our Pre-Strategy Intervention (PSI) is compared against the Base MARL, the GPSI and the Intrinsic Reward baselines.
To further validate the scalability of our Pre-Strategy Intervention, we conducted new experiments on the 4-player version of the Hanabi challenge. This setting serves as a more demanding benchmark for coordination, as the strategic complexity and communication burden are known to increase dramatically with the number of players [sudhakar2025generalist]. The results presented in Figure 14 demonstrate that our method achieves significantly higher scores than both the base MARL algorithms and the approach of the global intervention paradigm. This strong performance in a larger-scale setting provides robust evidence of our method’s ability to effectively handle more complex coordination challenges.
@@ I.7 Heterogeneous MPE
To evaluate the performance of our targeted intervention approach (Pre-Strategy Intervention) in settings with diverse agent capabilities, we designed a heterogeneous MPE scenario. This environment features three agents with asymmetric speeds: one “sprinter” agent with five times the normal speed and two normal-speed agents, with all methods built upon the IQL algorithm. The results, shown in Figure 15, reveal a consistent trend: intervening on a normal-speed agent leads to a noticeable improvement in final team performance (extrinsic return) compared to intervening on the uniquely capable sprinter, although the magnitude of this difference is modest in this specific environment. Our analysis suggests this is because the system’s primary limitation: the coordination bottleneck between the two normal-speed agents, when intervening on the sprinter agent. Even a subtle increase in the predictability of one normal agent’s behaviour through our intervention can alleviate this core challenge, leading to a more effective overall team strategy and demonstrating our approach’s potential in heterogeneous teams where targeted coordination adjustments can yield benefits.
I.8 Extending Pre-Strategy Intervention to Global Intervention Paradigm
(a) IPPO (extrinsic).
(b) IPPO (intrinsic).
(c) MAPPO (extrinsic).
(d) MAPPO (intrinsic).
(e) PQN-IQL (extrinsic).
(f) PQN-IQL (intrinsic).
(g) PQN-VDN (extrinsic).
(h) PQN-VDN (intrinsic).
Figure 16: Impact of intervention scope on performance in Hanabi. While extrinsic returns (task completion) are similar, Targeted Intervention (on a single agent) leads to higher per-agent intrinsic returns (convention adherence) compared to a Global Intervention baseline where all agents are guided towards the convention. Intrinsic returns are averaged per intervened agent.
To understand the impact of intervention scope, Figure 16 compares our standard Targeted Pre-strategy Intervention (PSI) with the Global Pre-Strategy Intervention variant (GPSI) where the pre-policy module is applied to guide all agents simultaneously towards the convention. While the performance on primary task completion (extrinsic returns) is broadly similar between these two intervention scopes across the tested MARL algorithms, a notable difference appears in the intrinsic rewards (e.g., as shown in subfigures like 16(b), 16(f), and 16(h) for specific backbones). When PSI is targeted to a single agent, that agent typically achieves a higher intrinsic return (representing convention adherence) than the average per-agent intrinsic return obtained by GPSI. This suggests that targeted intervention may lead to more focused and efficient convention learning by the targeted agent, potentially avoiding redundancies or conflicting convention-based actions that can arise when all agents are simultaneously guided by the same pre-strategy.
@@ I.8 Extending Pre-Strategy Intervention to Global Intervention Paradigm
Our work on Pre-Strategy Intervention (implementing the targeted intervention paradigm) in multi-agent reinforcement learning (MARL) introduces a theoretical framework for agent coordination. As foundational research conducted entirely in simulated environments, the immediate societal impact is limited. However, we acknowledge that theoretical advances often lay groundwork for future applications. If extended to practical domains, this research could potentially enable more efficient coordination in multi-agent systems by focusing guidance on specific agents rather than requiring direct control of an entire system. Our targeted intervention paradigm could reduce computational and communication overhead in domains like robotic teams, traffic management, or resource allocation. At the same time, we recognize that techniques for influencing multi-agent systems through targeted intervention could potentially be misapplied if deployed in socially sensitive contexts without appropriate oversight. Future work that moves this theoretical foundation toward real-world applications will carefully consider application-specific ethical implications and incorporate appropriate safeguards against potential misuse.